### v1

@@
111$\quad$1 School of Creative Media, City University of Hong Kong, Hong Kong, China. E-mail: Q. Wang, qiaowang-c@my.cityu.edu.hk; C. Xiao, chufeng.xiao@my.cityu.edu.hk; M. Lau, manfred.lau@cityu.edu.hk.
222$\quad$2 Division of Arts and Machine Creativity, Hong Kong University of Science and Technology, Hong Kong, China. E-mail: H. Fu, hongbofu@ust.hk.
@@ 1 Introduction
Understanding and simulating the human-object interaction (HOI) is a fundamental task for reconstructing realistic 3D worlds. Many learning-based methods have been proposed to address problems related to this topic, including 3D reconstruction [1, 2, 3, 4, 5, 6], object localization [7, 8], 3D object/hand tracking [9, 10, 11, 12, 13, 14, 15], and grasp generation [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. However, such methods need a large-scale and high-quality HOI dataset for training. The collection of such a dataset demands high-precision motion capture (MoCap) equipment and significant labor resources.
One possible way to acquire HOI data at a low cost is to produce a hand pose given the geometry of a target object. Another method (OakInk [26]) transfers a hand pose from existing HOI data (e.g., a hand pose grasping a source object scanned by MoCap) to a new object given the correspondence between the affordance maps (which indicates the contact points between a hand pose and an object) of the two objects. However, the method could not produce an accurate and plausible hand pose when the source object has a significant geometry difference from the target object, as shown in Figure 2a. This is because the transfer method highly relies on the whole shape of the two objects to compute the correspondences between them. Thus, the method struggles to transfer a hand pose between cross-category objects (e.g., mug and teapot), even though they have similar affordance in a local part.
@@
Based on the above observations, we introduce a novel part-based transfer method via the generalized cylinder (GC) representation to solve the problem of cross-category HOI transfer. Compared to the global-based transfer method, the part-based method can avoid the influence of shape differences in object regions that are irrelevant to the hand pose transfer. Additionally, using parts as the transferring units is more generic and allows for the use of parametric representations to represent contact regions. The generalized cylinder serves as a primitive representation of 3D object geometry, enabling the establishment of a unified and size-invariant correspondence between two different object parts. Specifically, we first model a generalized cylinder to abstract and parameterize the object parts of a source and target. Similar to the representation employed in GCD [27], a GC represents each point on the surface via the medial skeleton of an object part and cross-section profile curves along the skeleton. Following the polar coordinate systems created by the modeled GC, the contact points between a hand pose and a source part can be easily transferred to a target part. To maintain the target hand-grasping pose to be similar to the source pose, we estimate the target contact map while preserving the relative positions among contact points unchanged. Unlike methods that use registration or shape interpolation to find correspondences, the GC-based method mitigates the influence of size changes by preserving the relative positions among contact points. It also reduces the impact of shape differences by ensuring that contact points maintain their relative positions on the object globally. Finally, based on the re-targeted affordances, we refine a computed hand pose to approximately grasp the target object part through an optimization process, similar to the pose refinement method adopted in OakInk.
@@ 2 Related work
3D object decomposition decomposes complex objects into simple and semantic parts. It plays a pivotal role in the understanding and manipulation of 3D structures. For example, to train a learning-based method for hierarchical semantic segmentation, PartNet [28] presents a large-scale 3D object benchmark with fine-grained part-level annotations. Unlike learning-based methods, GCD proposes an optimization method to decompose a 3D object into cylindrical parts following the principle of geometry simplicity. However, the decomposition of GCD only considers the cylindricity measurement of object geometry but not the affordance involved by human hands. Different from the object decomposition task that decomposes a single object into parts, our work focuses on constructing the affordance correspondence between two object parts. To ensure the correspondence is size-invariant and unified, we utilize a generalized cylinder representation (similar to GCD) to represent an object part.
@@ 2.2 Object Affordance Estimation
Object affordance segments daily objects into different regions based on how humans grasp the objects. AffordanceNet and 3D affordanceNet [29, 30] propose learning-based methods trained on their prepared datasets to estimate object affordance given object shapes. Tactile Mesh Saliency [31] generates grasp saliency maps based on the geometry and visual features of 3D objects. However, such learning-based methods struggle to produce plausible affordance maps for out-of-domain data. To generate robust contact maps for optimizing a hand pose, we transfer a contact map from a source shape to a target shape, based on a unified representation of generalized cylinders.
@@ 2.3 Hand Grasp Synthesis
Many early works investigate grasp synthesis for robotics. Researchers use physics-based or deep learning methods to estimate grasping poses for robotic arms or grippers to grasp objects. GraspIt! [32] is an early open-source toolbox to generate grasps based on physical simulations of objects and grippers. SynGrasp [33] provides faster simulation in MATLAB Simulink. Some works [34, 35] use semantic recognition in grasping scenes to produce semantic-aware robot poses. ObMan [3] incorporates MANO [36], a parametric hand model, with GraspIt! to generate grasps based on the collected dataset of human hand images. Although such robotic-oriented methods can be applied to human hand grasp synthesis by mapping poses from grippers to human fingers, the methods would often produce unnatural poses for human grasping due to the gap between humans and robots.
With the advancement of deep learning and reinforcement learning techniques recently, researchers aim to develop a data-driven model to generate grasps based on the geometry of 3D objects. For example, ObMan trains a deep neural network on their proposed dataset to jointly reconstruct a hand pose and a 3D object model given a monocular RGB image. Many works propose diverse representations of HOIs and adopt cVAE-like models [37, 38] to generate such representations based on object geometry features. Specifically, GrabNet [18] and GraspTTA [16] trains cVAE networks with well-captured datasets to sample hand poses or contact maps given the geometry features of 3D shapes. Contact2Grasp [39] presents a novel hand pose refinement method that divides a hand into multiple regions and separately optimizes the hand regions to approximate a given contact map. Yang et al. [40] use anchors to represent the local regions of a hand and project the closest anchors from the hand to the object surface to form a contact map. Contactgen [23] introduces contact direction between a hand and an object surface to represent HOI data. Instead of generating static hand poses, ManipNet [41] generates a sequence of hand motions of interactions with objects based on hand wrist and object movements. Although these data-driven models can generate plausible results, they require high-quality data as input and lack controllability for hand pose outputs. In contrast, our method can produce a hand pose similar to a given pose to grasp a new object.
Reinforcement learning has been widely applied to grasp generation with the help of physical simulation engines. For example, She et al. [42] generate high degrees of freedom (DOFs) grasping poses between robot grippers and objects via reinforcement learning, while D-Grasp [25] synthesizes human hand motions. Although these data-driven methods can produce impressive hand poses, they highly rely on the training data and could not perform well for out-of-domain objects.
@@ 2.4 HOI data transferring
Transferring HOI data from source to target is a solution that can take advantage of high-fidelity HOI data from an existing small-scale dataset. OakInk transfers a contact map [40] from a source object to a target object through shape interpolation by DeepSDF [43]. However, as discussed in Section 1, this method suffers greatly from the issues caused by size and shape variance between source and target objects. TOCH [44] can transfer interaction sequences between objects with similar geometry features by learning to refine HOI sequences via GRU modules [45]. Similarly, we propose a transfer method to produce high-quality hand poses for grasping general objects. We compare our method with the existing transfer baseline (OakInk) to show its effectiveness and superiority (Section 4).
@@ 3.1 Generalized Cylinder Representation
When we use our hands to grasp a specific, semantically meaningful part of an object, our grasp pose tends to align with the primary structure or overall shape of that part. Minor structural details or small features have less influence on the grasp, as the hand primarily adapts to the core shape needed for a stable hold. As shown in Figure 4, some object parts are directly representable by generalized cylinders (Figure 4a), while other parts cannot be directly represented by generalized cylinders due to their complexity (Figure 4b). In our work, geometric abstraction is applied to these object parts for approximation by generalized cylinders. This approach approximates the object’s form by reducing complex features, such as intricate details, irregularities, or fine-scale variations, into simpler geometric structures. Instead of capturing all geometric details, we approximate the shape by filling holes, smoothing contours, or simplifying features, making the representation computationally efficient while still maintaining the essential structure of the part for hand grasping. With this geometric abstraction, the resulting transferred contact map generally enables reasonable grasp poses. The GC parameters are based on the GC representation of object parts. Specifically, we divide the skeleton of an object part into $n$+1 regions by evenly sampling $n$ cross-section curves to form a GC. For each cross-section curve, we represent it as a series of connected line segments that form a closed loop. Empirically, we found that $n$=30 is enough to maintain the necessary geometry of an object for estimating a target hand grasp. A larger $n$ leads to limited improvement but much lower computation efficiency.
@@ 3.2 GC-based Contact Transfer
A contact map consisting of several contact points between hand joints and object surface can effectively reflect how a hand grasps an object in the HOI data. Thus, we set up a contact map [40] as a medium to bridge the correspondence between two partial shapes. Unlike OakInk-based methods that are sensitive to object shapes and sizes due to shape-interpolated correspondences (Figure 1), we construct size-invariant correspondences via a GC representation to transfer a contact map between two GCs. We mainly describe below how to transfer the contact map from the source GC to the target GC. First, we transfer the contact map from a source shape to the modeled source GC by finding the closest point on the GC surface for each contact point.
@@ 4 Evaluation
We have conducted extensive experiments to evaluate PartHOI quantitatively and qualitatively. We first introduce the dataset and metrics we chose to test our method and baselines in Section 4.1 and Section 4.2. We establish three hand pose transfer baselines, i.e., vanilla OakInk based on the global shape of objects, part-based OakInk (P-OakInk) that directly applies OakInk to object parts, and direct shape-based hand pose transfer (P-reg) for object parts. For “P-reg”, we begin by applying shape-based fast global registration [46] to estimate an initial transformation matrix between source and target object parts. We then refine this alignment using Iterative Closest Point (ICP) [47], which iteratively minimizes the distance between corresponding points to register the contact map from the source to the target object. Finally, we apply a similar optimization approach as our method to fine-tune the hand pose based on the refined contact map. We show the comparisons between our method and the three baselines respectively for intra-category transfer (Section 4.3) and cross-category transfer (Section 4.4).
@@ 4.2 Metrics
Inspired by prior works [40, 16, 3, 26], we utilize four metrics to evaluate the synthesized HOI data, including penetration depth, penetration volume, disjointed distance, and simulation displacement.
@@
Penetration Volume (PV). PV quantifies the volume of intersection between a hand pose and the grasping object. To calculate PV, we first voxelize the object shape using a voxel grid size of 128 via voxelization methods [48, 49] and then count the number of hand-object intersected voxels.
@@ 4.3 Intra-category Transfer
PD $cm$downarrow 0.183 0.173 0.189 0.184
PV $cm^{3}$downarrow 1.423 1.191 1.177 1.365
DD $cm$downarrow 0.963 1.080 1.075 1.025
[…]
SD $cm$downarrow 2.581 2.613 2.712 2.717
@@ 4.4 Cross-category Transfer
PD $cm$downarrow 0.195 0.334 0.197 0.312
PV $cm^{3}$downarrow 1.619 4.158 1.461 3.740
DD $cm$downarrow 1.315 3.796 1.530 1.752
[…]
SD $cm$downarrow 3.051 3.421 3.193 3.250
@@ 5 Conclusions and Discussion
In this work, we present PartHOI, a part-based HOI transfer method, to produce plausible hand poses to grasp new objects given source HOI data. Based on the constructed GC-based correspondence, our method can transfer contact maps between two object parts with high shape and size differences. The extensive evaluation demonstrates that our method performs well in HOI transfer and is superior to the existing methods, especially for cross-category transfer. It implies that our method can effectively expand the scale of existing HOI datasets by propagating the existing hand poses to novel categories of objects. While our method works well for HOI transfer, there are some limitations. First, we manually segment grasping parts and extract their skeletons for GC modeling since existing automatic methods [50, 51, 28] could not produce satisfactory results under diverse grasping intentions. In the future, we are interested in adapting such methods to our task. Second, our method may struggle to handle objects with highly complicated structures that GC could not model well. The abstraction of object parts tends to overlook minor structures, which can lead to inaccuracies when hands interact with these finer details. For example, our method does not handle scenarios where fingers interact with small-scale object features, such as inserting into tiny holes or pressing small buttons on surfaces. In the future, we can incorporate geometric descriptors (e.g., SIFT/SURF features) that capture local surface variations with a GC representation to build a more comprehensive correspondence. Third, as shown in Figure 12, our method can produce results with interpenetration when dealing with object parts that have very thin structures, such as the eyeglasses’ frames. This issue arises because the signed distance values for these surfaces are too small to effectively prevent penetration during the optimization process. In future work, we plan to address this limitation by assigning greater weights to the signed distances of thin surfaces based on the GC parameters, thereby reducing interpenetration in the optimization step.

### v2

@@
$1\qquad$ School of Creative Media, City University of Hong Kong, Hong Kong, China. E-mail: Q. Wang, qiaowang-c@my.cityu.edu.hk; C. Xiao, chufeng.xiao@my.cityu.edu.hk.
$2\qquad$ Department of Computer Science, Lakehead University, Orillia, Ontario, Canada. E-mail: M. Lau, manfred.lau@lakeheadu.ca.
$3\qquad$ Division of Arts and Machine Creativity, Hong Kong University of Science and Technology, Hong Kong, China. E-mail: H. Fu, hongbofu@ust.hk.
@@ 1 Introduction
Understanding and simulating the human-object interaction (HOI) is a fundamental task for reconstructing realistic 3D worlds. Many learning-based methods have been proposed to address problems related to this topic, including 3D reconstruction [cao_reconstructing_2021, hasson_towards_2021, hasson_learning_2019, tzionas_3d_2015, chen_gsdf:_2023, li_chord:_2023], object localization [chalon_online_2013, Lee_2020_WACV], 3D object/hand tracking [Tsoli_2018_ECCV, hamer_tracking_2009, leibe_real-time_2016, krainin_manipulator_2011, rangesh_hidden_2016, mueller_real-time_2017, panteleris_back_2017], and grasp generation [jiang_hand-object_2021, Brahmbhatt_2020_ECCV, GRAB:2020, brahmbhatt_contactgrasp:_2019, grady_contactopt:_2021, zhu_toward_2021, xu_unidexgrasp:_2023, liu_contactgen:_2023, brahmbhatt_contactdb:_2019, christen2022d]. However, such methods need a large-scale and high-quality HOI dataset for training. The collection of such a dataset demands high-precision motion capture (MoCap) equipment and significant labor resources.
One possible way to acquire HOI data at a low cost is to produce a hand pose given the geometry of a target object. Another method (OakInk [yang_oakink:_2022]) transfers a hand pose from existing HOI data (e.g., a hand pose grasping a source object scanned by MoCap) to a new object given the correspondence between the affordance maps (which indicates the contact points between a hand pose and an object) of the two objects. However, the method could not produce an accurate and plausible hand pose when the source object has a significant geometry difference from the target object, as shown in Figure 2a. This is because the transfer method highly relies on the whole shape of the two objects to compute the correspondences between them. Thus, the method struggles to transfer a hand pose between cross-category objects (e.g., mug and teapot), even though they have similar affordance in a local part.
@@
Based on the above observations, we introduce a novel part-based transfer method via the generalized cylinder (GC) representation to solve the problem of cross-category HOI transfer. Compared to the global-based transfer method, the part-based method can avoid the influence of shape differences in object regions that are irrelevant to the hand pose transfer. Additionally, using parts as the transferring units is more generic and allows for the use of parametric representations to represent contact regions. The generalized cylinder serves as a primitive representation of 3D object geometry, enabling the establishment of a unified and size-invariant correspondence between two different object parts. Specifically, we first model a generalized cylinder to abstract and parameterize the object parts of a source and target. Similar to the representation employed in GCD [gcd15], a GC represents each point on the surface via the medial skeleton of an object part and cross-section profile curves along the skeleton. Following the polar coordinate systems created by the modeled GC, the contact points between a hand pose and a source part can be easily transferred to a target part. To maintain the target hand-grasping pose to be similar to the source pose, we estimate the target contact map while preserving the relative positions among contact points unchanged. Unlike methods that use registration or shape interpolation to find correspondences, the GC-based method mitigates the influence of size changes by preserving the relative positions among contact points. It also reduces the impact of shape differences by ensuring that contact points maintain their relative positions on the object globally. Finally, based on the re-targeted affordances, we refine a computed hand pose to approximately grasp the target object part through an optimization process, similar to the pose refinement method adopted in OakInk.
@@ 2 Related work
3D object decomposition decomposes complex objects into simple and semantic parts. It plays a pivotal role in the understanding and manipulation of 3D structures. For example, to train a learning-based method for hierarchical semantic segmentation, PartNet [mo2019partnet] presents a large-scale 3D object benchmark with fine-grained part-level annotations. Unlike learning-based methods, GCD proposes an optimization method to decompose a 3D object into cylindrical parts following the principle of geometry simplicity. However, the decomposition of GCD only considers the cylindricity measurement of object geometry but not the affordance involved by human hands. Different from the object decomposition task that decomposes a single object into parts, our work focuses on constructing the affordance correspondence between two object parts. To ensure the correspondence is size-invariant and unified, we utilize a generalized cylinder representation (similar to GCD) to represent an object part.
@@ 2.2 Object Affordance Estimation
Object affordance segments daily objects into different regions based on how humans grasp the objects. AffordanceNet and 3D affordanceNet [do_affordancenet:_2018, deng_3d_2021] propose learning-based methods trained on their prepared datasets to estimate object affordance given object shapes. Tactile Mesh Saliency [Lau_tactile] generates grasp saliency maps based on the geometry and visual features of 3D objects. However, such learning-based methods struggle to produce plausible affordance maps for out-of-domain data. To generate robust contact maps for optimizing a hand pose, we transfer a contact map from a source shape to a target shape, based on a unified representation of generalized cylinders.
@@ 2.3 Hand Grasp Synthesis
Many early works investigate grasp synthesis for robotics. Researchers use physics-based or deep learning methods to estimate grasping poses for robotic arms or grippers to grasp objects. GraspIt! [miller_graspit!_2004] is an early open-source toolbox to generate grasps based on physical simulations of objects and grippers. SynGrasp [6630708] provides faster simulation in MATLAB Simulink. Some works [dang_semantic_2012, detry_task-oriented_2017] use semantic recognition in grasping scenes to produce semantic-aware robot poses. ObMan [hasson_learning_2019] incorporates MANO [MANO:SIGGRAPHASIA:2017], a parametric hand model, with GraspIt! to generate grasps based on the collected dataset of human hand images. Although such robotic-oriented methods can be applied to human hand grasp synthesis by mapping poses from grippers to human fingers, the methods would often produce unnatural poses for human grasping due to the gap between humans and robots.
With the advancement of deep learning and reinforcement learning techniques recently, researchers aim to develop a data-driven model to generate grasps based on the geometry of 3D objects. For example, ObMan trains a deep neural network on their proposed dataset to jointly reconstruct a hand pose and a 3D object model given a monocular RGB image. Many works propose diverse representations of HOIs and adopt cVAE-like models [kingma2013auto, rezende2014stochastic] to generate such representations based on object geometry features. Specifically, GrabNet [GRAB:2020] and GraspTTA [jiang_hand-object_2021] trains cVAE networks with well-captured datasets to sample hand poses or contact maps given the geometry features of 3D shapes. Contact2Grasp [ijcai2023p117] presents a novel hand pose refinement method that divides a hand into multiple regions and separately optimizes the hand regions to approximate a given contact map. Yang et al. [yang_cpf:_2021] use anchors to represent the local regions of a hand and project the closest anchors from the hand to the object surface to form a contact map. Contactgen [liu_contactgen:_2023] introduces contact direction between a hand and an object surface to represent HOI data. Instead of generating static hand poses, ManipNet [zhang_manipnet:_2021] generates a sequence of hand motions of interactions with objects based on hand wrist and object movements. Although these data-driven models can generate plausible results, they require high-quality data as input and lack controllability for hand pose outputs. In contrast, our method can produce a hand pose similar to a given pose to grasp a new object.
Reinforcement learning has been widely applied to grasp generation with the help of physical simulation engines. For example, She et al. [she_learning_2022] generate high degrees of freedom (DOFs) grasping poses between robot grippers and objects via reinforcement learning, while D-Grasp [christen2022d] synthesizes human hand motions. Although these data-driven methods can produce impressive hand poses, they highly rely on the training data and could not perform well for out-of-domain objects.
@@ 2.4 HOI data transferring
Transferring HOI data from source to target is a solution that can take advantage of high-fidelity HOI data from an existing small-scale dataset. OakInk transfers a contact map [yang_cpf:_2021] from a source object to a target object through shape interpolation by DeepSDF [park_deepsdf:_2019]. However, as discussed in Section 1, this method suffers greatly from the issues caused by size and shape variance between source and target objects. TOCH [zhou2022toch] can transfer interaction sequences between objects with similar geometry features by learning to refine HOI sequences via GRU modules [chung2014empirical]. Similarly, we propose a transfer method to produce high-quality hand poses for grasping general objects. We compare our method with the existing transfer baseline (OakInk) to show its effectiveness and superiority (Section 4).
@@ 3.1 Generalized Cylinder Representation
When we use our hands to grasp a specific, semantically meaningful part of an object, our grasp pose tends to align with the primary structure or overall shape of that part. Minor structural details or small features have less influence on the grasp, as the hand primarily adapts to the core shape needed for a stable hold. As shown in Figure 4, some object parts are directly representable by generalized cylinders (Figure 4a), while other parts cannot be directly represented by generalized cylinders due to their complexity (Figure 4b). In our work, geometric abstraction is applied to these object parts for approximation by generalized cylinders. This approach approximates the object’s form by reducing complex features, such as intricate details, irregularities, or fine-scale variations, into simpler geometric structures. Instead of capturing all geometric details, we approximate the shape by filling holes, smoothing contours, or simplifying features, making the representation computationally efficient while still maintaining the essential structure of the part for hand grasping. With this geometric abstraction, the resulting transferred contact map generally enables reasonable grasp poses. The GC parameters are based on the GC representation of object parts. Specifically, we divide the skeleton of an object part into nn+1 regions by evenly sampling $n$ cross-section curves to form a GC. For each cross-section curve, we represent it as a series of connected line segments that form a closed loop. Empirically, we found that nn=30 is enough to maintain the necessary geometry of an object for estimating a target hand grasp. A larger $n$ leads to limited improvement but much lower computation efficiency.
@@ 3.2 GC-based Contact Transfer
A contact map consisting of several contact points between hand joints and object surface can effectively reflect how a hand grasps an object in the HOI data. Thus, we set up a contact map [yang_cpf:_2021] as a medium to bridge the correspondence between two partial shapes. Unlike OakInk-based methods that are sensitive to object shapes and sizes due to shape-interpolated correspondences (Figure 1), we construct size-invariant correspondences via a GC representation to transfer a contact map between two GCs. We mainly describe below how to transfer the contact map from the source GC to the target GC. First, we transfer the contact map from a source shape to the modeled source GC by finding the closest point on the GC surface for each contact point.
@@ 4 Evaluation
We have conducted extensive experiments to evaluate PartHOI quantitatively and qualitatively. We first introduce the dataset and metrics we chose to test our method and baselines in Section 4.1 and Section 4.2. We establish three hand pose transfer baselines, i.e., vanilla OakInk based on the global shape of objects, part-based OakInk (P-OakInk) that directly applies OakInk to object parts, and direct shape-based hand pose transfer (P-reg) for object parts. For “P-reg”, we begin by applying shape-based fast global registration [zhou2016fast] to estimate an initial transformation matrix between source and target object parts. We then refine this alignment using Iterative Closest Point (ICP) [Rusinkiewicz2001Iterative], which iteratively minimizes the distance between corresponding points to register the contact map from the source to the target object. Finally, we apply a similar optimization approach as our method to fine-tune the hand pose based on the refined contact map. We show the comparisons between our method and the three baselines respectively for intra-category transfer (Section 4.3) and cross-category transfer (Section 4.4).
@@ 4.2 Metrics
Inspired by prior works [yang_cpf:_2021, jiang_hand-object_2021, hasson_learning_2019, yang_oakink:_2022], we utilize four metrics to evaluate the synthesized HOI data, including penetration depth, penetration volume, disjointed distance, and simulation displacement.
@@
Penetration Volume (PV). PV quantifies the volume of intersection between a hand pose and the grasping object. To calculate PV, we first voxelize the object shape using a voxel grid size of 128 via voxelization methods [binvox, nooruddin03] and then count the number of hand-object intersected voxels.
@@ 4.3 Intra-category Transfer
PD cmcm $\downarrow$ 0.183 0.173 0.189 0.184
PV $cm^{3}$ $\downarrow$ 1.423 1.191 1.177 1.365
$D$ cmcm $\downarrow$ 0.963 1.080 1.075 1.025
CR $\%$ $\uparrow$ 24.94 24.34 24.68 24.26
SD cmcm $\downarrow$ 2.581 2.613 2.712 2.717
@@ 4.4 Cross-category Transfer
PD cmcm $\downarrow$ 0.195 0.334 0.197 0.312
PV $cm^{3}$ $\downarrow$ 1.619 4.158 1.461 3.740
$D$ cmcm $\downarrow$ 1.315 3.796 1.530 1.752
CR $\%$ $\uparrow$ 28.80 21.79 27.96 27.89
SD cmcm $\downarrow$ 3.051 3.421 3.193 3.250
@@ 5 Conclusions and Discussion
In this work, we present PartHOI, a part-based HOI transfer method, to produce plausible hand poses to grasp new objects given source HOI data. Based on the constructed GC-based correspondence, our method can transfer contact maps between two object parts with high shape and size differences. The extensive evaluation demonstrates that our method performs well in HOI transfer and is superior to the existing methods, especially for cross-category transfer. It implies that our method can effectively expand the scale of existing HOI datasets by propagating the existing hand poses to novel categories of objects. While our method works well for HOI transfer, there are some limitations. First, we manually segment grasping parts and extract their skeletons for GC modeling since existing automatic methods [rosa_sig09, Huang2013, mo2019partnet] could not produce satisfactory results under diverse grasping intentions. In the future, we are interested in adapting such methods to our task. Second, our method may struggle to handle objects with highly complicated structures that GC could not model well. The abstraction of object parts tends to overlook minor structures, which can lead to inaccuracies when hands interact with these finer details. For example, our method does not handle scenarios where fingers interact with small-scale object features, such as inserting into tiny holes or pressing small buttons on surfaces. In the future, we can incorporate geometric descriptors (e.g., SIFT/SURF features) that capture local surface variations with a GC representation to build a more comprehensive correspondence. Third, as shown in Figure 12, our method can produce results with interpenetration when dealing with object parts that have very thin structures, such as the eyeglasses’ frames. This issue arises because the signed distance values for these surfaces are too small to effectively prevent penetration during the optimization process. In future work, we plan to address this limitation by assigning greater weights to the signed distances of thin surfaces based on the GC parameters, thereby reducing interpenetration in the optimization step.